An institute reported that 63​% of its members indicate that lack of ethical culture within financial firms has contributed most to the lack of trust in the financial industry. Suppose that you select a sample of 100 institute members. Complete parts ​(a) through ​(d) below.

B. The probability is 70​% that the sample percentage will be contained within what symmetrical limits of the population​ percentage?

The probability is 70​% that the sample percentage will be contained above ____% and below ____​%.
Answer #1

solution

p = 0.63 , q = 0.37, n = 100

Mean = p = 0.63

standard deviation = sqrt(npq) = 0.0483

for middle 70%

z = 1.04

lower value = 0.63-1.04*0.0483 = 0.5798

upper value = 0.63+1.04*0.0483 = 0.6802

The probability is 70% that the sample percentage will be contained above 57.98 % and below 68.02%